OpenStudy (anonymous):

Evaluate each expression for u = -5 and v = 3. 2u - 8 + 3v A) 1 7v - u + 4 B) 30 11 - 3v - u C) -9 9 + u - v D) 7

OpenStudy (brooke2353):

C.

OpenStudy (anonymous):

I need all of them

OpenStudy (anonymous):

In order please

OpenStudy (brooke2353):

those*

OpenStudy (anonymous):

@brooke2353 can you help. Or explain how

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

What? You have to match ABC and D with the problems

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@IMstuck

OpenStudy (mathstudent55):

Take each expression and replace u with -5 and v with 3. Then evaluate the expression. Here is the first expression: 2u - 8 + 3v First, substitute the values for u and v: 2(-5) - 8 + 3(3) Now evaluate that expression: = -10 - 8 + 9 = -18 + 9 = -9 The answer to the first expression is C. -9 Now do the same for the other expressions.

OpenStudy (anonymous):

Wow thanks your a lifesaver

OpenStudy (mathstudent55):

You're welcome.